OpenStudy (anonymous):

savings account compounds interest, at a rate of 17%, once a year. John puts $1,000 in the account as the principal. How can John set up a function to track the amount of money he has?

OpenStudy (anonymous):

Compound intrest formula: \[A = P (1+ \frac{ r }{ n })nt\] A = Amount at maturity or at time t. P = Principal Amount = $1,000 r = Annual interest rate in decimal = 0.17 n = compounding period (compounded how many times a year) = 1 t = years invested

OpenStudy (anonymous):

so i would use this equation to solve the problem? @Light&Happiness

OpenStudy (anonymous):

@Light&Happiness

OpenStudy (anonymous):

Yes, use the equation given! :)

OpenStudy (anonymous):

Wait what i put for T @Light&Happiness

OpenStudy (anonymous):

T = years invested

OpenStudy (anonymous):

You can use http://www.wolframalpha.com/ to help you solve this problem.

OpenStudy (anonymous):

I think responder meant to write A = P(1+r/n)^nt